OpenStudy (anonymous):

Solve. Round your answer to three decimal places 4^x+1=6^x-1

OpenStudy (anonymous):



OpenStudy (anonymous):

take the log of both sides\[\log \left( 4^{x+1} \right) = \log \left( 6^{x-1} \right)\]and then solve for x by using algebra and properties of logarithms

OpenStudy (anonymous):

well i got x = 1 then what do i do ?

OpenStudy (anonymous):

These are the answer choices.. A. .128 B..200 C. 7.838 D. 5.00

OpenStudy (anonymous):

how did you get 1 ?

OpenStudy (anonymous):

through the calculator :/

OpenStudy (anonymous):

recheck your work, the answer is C